OpenStudy (anonymous):

find the sum of the finite geometric series by using the formula for sn 1-1/2+1/4-1/8+1/16-1/32+1/64-1/128 the sum of the finite geometric series is???

OpenStudy (anonymous):

a=1/2 r=-1/2 S=a/ (1-r)

OpenStudy (anonymous):

\[Sn = (a(1-r^n)) / (1-r)\]

OpenStudy (anonymous):

=(1/2) / (3/2) = (1/2) (2/3) = (1/3)

OpenStudy (anonymous):

oh i thought it said infinite, but yeh same thing, it will be extremely close to 1/3

OpenStudy (anonymous):

thank u...